$SPAD/src/input d02bbf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
xend:SF:=8.0
n:=3
irelab:=0
x:SF:=0.0
y:Matrix SF:=
 [[0.0 ,0.5 ,%pi*0.2 ]]
tol:SF:=0.0001
vef:Vector Expression Float:=
 [tan(Y[3]) ,-0.032*tan(Y[3])/Y[2] -0.02*Y[2]/cos(Y[3]) ,-0.032/(Y[2]**2) ]
fcn:Asp7(FCN):= retract vef
vm:Vector MachineFloat:=
 [1,2,3,4,5,6,7,8]
output:Asp8(OUTPUT):= coerce vm
result:=d02bbf(xend,# vm,n,irelab,x,y,tol,-1,fcn,output)

———-

References^{[1] nothing}

4